OpenStudy (kiana1345):

For each function, (a) determine whether the graph opens upward or downward, (b) find the axis of symmetry, (c) find the vertex, (d) find the y-intercept, and (e) graph the function. g(x) = -3x^2 + 6x

OpenStudy (mathmale):

g(x) = -3x^2 + 6x is a quadratic function of the form y=ax^2+bx+c. Here, a=-3, b=6 and c=0. Because the first coefficient (a=-3) is negative, the graph of this quadratic, a parabola, opens DOWN. If you want to use a calculus approach, find the derivative of g(x) = -3x^2 + 6x, and then the 2nd derivative. You will find that the 2nd derivative is negative. If the 2nd deriv. of a function is neg., that means that the graph of the function at and near certain x values opens DOWN. Setting calculus aside for now, focus on the other questions. How would you find the axis of symmetry? Using the coefficients a=-3 and b=6, find x=-b/(2a). How would you then find the coordinates of the vertex? Hint: the x-coordinate is the same as that of the axis of symmetry. Please show all work on which you want commentary.

OpenStudy (mathmale):

A hearty welcome to OpenStudy!